# Two-Stage Stance Labeling: User-Hashtag Heuristics with Graph Neural Networks thanks: Army Research Office

### v1

@@ Abstract
The high volume and rapid evolution of content on social media present major challenges for studying the stance of social media users. In this work, we develop a two stage stance labeling method that utilizes the user-hashtag bipartite graph and the user-user interaction graph. In the first stage, a simple and efficient heuristic for stance labeling uses the user-hashtag bipartite graph to iteratively update the stance association of user and hashtag nodes via a label propagation mechanism. This set of soft labels is then integrated with the user-user interaction graph to train a graph neural network (GNN) model using semi-supervised learning. We evaluate this method on two large-scale datasets containing tweets related to climate change from June 2021 to June 2022 and gun control from January 2022 to January 2023. Experiments demonstrate that our user-hashtag heuristic and the semi-supervised GNN method outperform zero-shot stance labeling using LLMs such as GPT4. Further analysis illustrates how the stance labeling information and interaction graph can be used for evaluating the polarization of social media interactions on divisive issues such as climate change and gun control.
@@ I Introduction
Stance labeling remains a particularly challenging task due to the informal and noisy nature of social media language, the presence of sarcasm and irony, and the potential biases introduced by human annotators and automated models. To further facilitate research into polarization online, we develop a stance labeling method that incorporates both textual content and user social interactions in a scalable two-stage pipeline. In the first stage, we apply a reciprocal label propagation algorithm to the user-hashtag bipartite graph that generates a soft-labeled set of users. In the second stage, we construct a signed, weighted, and attributed user-user interaction network using transformer language models to embed the content of user posts and to determine the sentiment of interactions between users. Using the soft labels generated from the label propagation stage, we employ semi-supervised training to train a graph neural network classifier to classify user stances.
Our stance labeling approach utilizes an understanding about linguistic and network homophily. Linguistic homophily states that people who share similar ideologies also share similar textual content, not only through similar keyword and hashtag usages but also through sentiment and linguistic features [8, 9], and the inverse phenomenon-that users with dissimilar views use different language-has been investigated by researchers [2]. Likewise, the concept of network homophily, which describes the phenomenon that similar nodes tend to be connected to one another has long been studied in graph theory [10, 11] and informs the success of graph neural networks [12, 13, 14]. By combining the content-based and network-based approaches into a single, scalable pipeline for stance labeling, we facilitate further large-scale analysis of both interactional and affective polarization online.
@@
  • We develop an approach for estimating user stance labels that incorporates both textual features and social network interactions
  • A two-stage pipeline allows our approach to scale to large datasets with soft labeling and semi-supervised training
  • We evaluate our method on two large-scale datasets on two contentious political issues: climate change and gun control
@@ II Related Work
As social media has become the primary forum for discussion on many topics, including political issues, there has been growing interest in user stance detection . Many existing methods can be divided into two broad categories: textual-based methods that use the content of messages, and network-based methods that use social or conversational structure. Textual-based methods use the content of tweets, user profiles, and hashtags to determine a user’s stance on a particular topic. Several studies have used supervised classification of user stance labels. Classifiers are trained using a set of features encoding tweet text, hashtags, and user profile information. These techniques have been applied to classify user political ideology . These approaches are generally rely on manually annotated sets of users to train the classifier models, and such methods are often limited by the size and diversity of the training content as manual annotation is both time-consuming and expensive. Furthermore, these methods can be susceptible to annotator bias poor generalizability due to the limited distribution of training examples.
@@
Topic modeling and user clustering are two unsupervised techniques that have been employed for stance labeling. These unsupervised textual generally embed tweets or user profile information in a latent feature space and then apply a variety of topic modeling or cluster methods to identify tweets of similar topics or clusters of similar users. While such techniques avoid the issues with manual annotation for supervised training, they suffer from the  curse of dimensionality—the search space for a solution grows exponentially with the increase in dimensionality as there are many more possible patterns than in lower-dimensional subspaces. Selecting useful and informative features, and the computation and memory complexity of clustering methods on high-dimensional create problems for unsupervised clustering-based methods. For high-dimensional data, many clustering techniques fail to produce meaningful clusters, but on the other hand, such techniques can be very efficient on low-dimensional features spaces . Several studies have therefore employed dimensionality reduction techniques prior to clustering to first project tweets to a low-dimensional space, and such techniques can be robust to class imbalance between different stance groups.
@@
Network-based approaches use the structure of social interaction networks to determine user ideology or stance. Such methods leverage techniques such as label propagation and min-cut, and node embedding techniques to leverage structural information from retweet  and user follower  networks in order to train. Xiao et al.  generated a multi-relational network using retweets, mentions, likes, and follows for binary classification of user ideologies.
@@
More recent works have combined both textual content and network features in socially-infused text mining. Li and Goldwasser  combine user interactions and user sharing of news media to detect bias in news articles, and Johnson, Jin, and Goldwasser  combine multiple features from lexical features and social interactions to categorize tweets by politicians on a varity of political topics. Pan et al.  use the network structure and node content to learn node representations for classifying the category of scientific publications. Such works have demonstrated improvements over earlier methods that rely solely on textual content or network structure alone.
@@
Our approach is similar in its reliance on socially-infused text mining for stance labeling, but our approach is unique from the methods described above in that we i) combine language features from state-of-the-art transformer language models with the structure and sentiment of social interactions using graph neural networks, and ii) our two-stage process allows our method to scale to large datasets with minimal supervision.
@@ III Methods
This section describes our proposed method to estimate the stance of users as a binary classification problem. We first describe the collection of data for the gun control and climate changes datasets. Then we characterize the hashtag-based heuristic method to generate soft labels for two stance groups of users, which are used as seed users for training and evaluating stance labeling methods. We then introduce several baseline methods followed by our GNN-based approach for stance labeling.
@@ III-A Data Collection
We collect tweets related to two prominent and contentious political issues—climate change and gun control. The keywords listed in Table  I are used to scrape the initial set of relevant tweets related to each topic. To fully capture conversation cascades and user interactions, we then recursively collect all referenced tweets from each tweet in the initially matched set. This ensures that all available tweets from relevant conversations are included in our analysis, even if particular tweets do not explicitly use any of the keywords. The climate change dataset consists of tweets published between June 1st, 2021 and May 31st, 2022. In all, the climate change dataset consists of a total of 46M tweets authored by 4.8M unique users and contains conversation threads of at least three tweets. The gun control dataset consists of tweets published between January 1st, 2022 and December 31st, 2022. In all, the gun control dataset dataset consists of a total of 14.4M tweets from 2.66M unique users containing 335,000 conversation threads of at least three tweets.
@@
Not only are these two topics highly relevant political issues that generate a large amount of engagement, news coverage, and user interaction on social media, but they are also good models for understanding the differences in behavior on social media for topics that are highly event driven, like gun control, and those that are less focused on particular events, like climate change. Figure 1 illustrates that conversation threads related to climate change are started at a fairly consistent rate over time. By contrast, Figure 2 shows that the frequency of discussions surrounding gun control spikes sharply in response to mass shooting events. The summer of 2022 saw several deadly and highly publicized mass shootings in Buffalo, Uvalde, and Chicago that generated large amounts of discourse surrounding gun control and gun violence on social media, illustrated by the steep slope from the end of May through late July. Shortly after mass shooting events though, the level of discourse abates, indicated by a flatter slope between early August until November, when a shooting event at UVA caused another spike in conversations. Event driven issues like gun control exhibit different posting behavior and user interactions that we hypothesize leads to differences in polarization and likely would require different techniques for intervening and mitigating polarization. Developing approaches for stance labeling that are both topic independent and robust to different types of discourse online is an important step for facilitating further analysis of polarization online.
Figure 1: Gantt plot of the largest Twitter conversation threads related to climate change.
Figure 2: Gantt plot of the largest Twitter conversation threads related to gun control.
@@ III-B User-Hashtag Stance Labeling
Based on a small (3-5) set of seed hashtags associated with each stance, we first propagate the stance associated with each seed hashtag across the user-hashtag bipartite graph to the set of users who posted a given labeled hashtag. The importance of a hashtag’s stance to a particular destination user is weighted by the number of time the user used each hashtag, as described in Algorithm 2. Users are then assigned a label based on the number of hashtag usages for each stance group. A user will be assigned a stance label based on the count of usages of labeled hashtags for each stance group. User labels are then propagated back to the hashtags used by each user, as described in Algorithm 3. Each hashtag is then scored based on the difference between the normalized count of usages from users of each stance group. Hashtag scores are normalized via min-max normalization and the mean and standard deviation of hashtag scores is computed. Hashtags that score one or more standard deviation above or below the mean are assigned the corresponding stance label. We then repeat alternating label propagation until the model converges, i.e. the set of labeled users is the same from the prior iteration to the next.
@@ III-C GNN Stance Labeling
While the user-hashtag stance labeling method can efficiently determine the stance group of users and discover important hashtags associated with each stance group, it is limited by the fact that not all users post hashtags and that hashtags consist of a small part of the overall textual content contained in tweets. Furthermore, the method does not take into account any social interactions between users. The method has high precision but relatively low recall due to the limited breadth of users that stance labeling method, we employ the user-hashtag bipartite method to generate a soft-labeled set of users from which we train a graph neural network classifier over the attributed user-user interaction graph.
We define a user interaction as a retweet, mention, reply, or quote from the author of the tweet towards another user. We then construct a weighted, signed, and attributed user-user interaction network $G=(V,E,X)$, where $V$ is the set of users and a directed edge $(u,v)_{i}\in E$ connects the author user to the target user of their interaction as defined above. The edge weight $w(u,v)_{i}$ corresponds to the sentiment of the tweet, scored by a BERTweet-based sentiment analysis model. BERTweet is a RoBERTa model fine-tuned on English tweets and trained on the SemEval 2017 corpus. To characterized the overall interactions between users, we consolidated the multiple edges between pairs of users corresponding to individual tweets by combining each directed edge $(u,v)_{i}$ into a single edge for each user pair $(u,v)$ with the composite edge weight $w(u,v)$ computed as the mean sentiment score across all tweets between two users. Each user $u\in V$ is also associated with a feature vector $x_{u}\in X$ which corresponds to the element-wise mean of the BERTweet embedding for each tweet authored by the user. We then train a graph neural network classifier over the user interaction graph, using the set of seed users labeled by the hashtag method as a training set.
We consider three implementations from two prominent families of graph neural networks convolutional and attentional. The neighbor coefficients may be fixed weights as in convolutional GNNs or may be computed implicitly, as in attentional GNNs. Specifically, we evaluate graphSAGE[13] and Graph Isomorphism Network (GIN) as examples of convolutional GNNs, and graph attention network (GAT) as an example attentional GNN.
@@
where $\mathcal{\tilde{N}}(v)$ is the local neighborhood of node $v$ with added self-loop, $\boldsymbol{\mathrm{W}}$ is a trainable weight matrix, $\sigma$ is a non-linear activation function. We use the mean aggregator to aggregate the neighboring node features.
Another convolutional GNN we consider is GIN. The message passing function in GIN is defined as:
$\boldsymbol{\hat{h}}_{v}^{(k)}=\mathrm{MLP}\left((1+\epsilon)\cdot\boldsymbol{h}_{v}^{(k-1)}+\sum_{u\in\mathcal{N}(v)}\boldsymbol{h}_{u}^{(k-1)}\right)$ (2)
Compared with the graphSAGE operator, GIN notably replaces the single layer transformation matrix with a multi-layer perceptron (MLP)[popescu2009multilayer] and uses a simple unnormalized summation aggregator over the local neighboring nodes rather than a mean aggregator.
Thirdly, we also consider a GAT layer operator, defined as:
@@
III-D Heuristic Label Validation
The hashtag-based heuristic is limited in its ability to label all users, covering only ()percent of the users in the gun control dataset and ()of users in the climate change dataset. Despite coverage limitations, the method works due well to its simplicity and speed for extracting an initial set of labeled users that can serve as a training set for other language-based and network-based stance labeling methods methods. A social science expert independently annotated a set of 250 and 350 users for each dataset by analyzing the top 20 most popular tweets from each user.
@@ III-D Heuristic Label Validation
GIN
III-E Baselines
We compare our stance labeling method against three transformer models that leverage the language content of users’ posts. We compute an embedding for each user by computing the element-wise mean of the embedding of all of the users’ tweets from BERT, RoBERTa, and BERTweet. We then train a classifier MLP head using the labels from the hashtag heuristic to predict user stance label. We evaluate the performance of both the transformer and GNN models using 5-fold cross validation on the large set of labeled users from the reciprocal stance labeling heuristic.
We further evaluate classification performance of stance labeling methods on a small set of users manually annotated by domain experts. Lastly, we compare the performance of state-of-the-art large language models (LLMs) against the aforementioned transformer-based language models and GNNs using zero-shot classification [15, 16] with GPT4 to predict user stances for the same set of users annotated manually by human experts.
We use function calling to prompt GPT-4 to provide the sentiment of content with respect to gun control or climate change. For the gun control dataset, the function call prompt is: ”The stance of the content. Is the message pro gun control, anti gun control, or neutral¿‘ with three classification options of ”pro”, ”anti”, or ”neutral”. For the climate change dataset, ”the stance of the content. Does the message capture the user’s belief in climate change, disbelief, or is it just neutral?” with classification options of ”belief”, ”disbelief”, or ”neutral”. We then aggregate the stance labeling results for each individual tweet to determine a user’s overall stance on the issue of gun control or climate change. Users who had a majority of tweets classified by one stance were labeling accordingly and those users whose tweets were labeled as all neutral are classified as undetermined. In the case where there is an equivalent number of tweets labeled as both stances, the user is labeled as undetermined.

### v2

@@ Abstract
The high volume and rapid evolution of content on social media present major challenges for studying the stance of social media users. In this work, we develop a two stage stance labeling method that utilizes the user-hashtag bipartite graph and the user-user interaction graph. In the first stage, a simple and efficient heuristic for stance labeling uses the user-hashtag bipartite graph to iteratively update the stance association of user and hashtag nodes via a label propagation mechanism. This set of soft labels is then integrated with the user-user interaction graph to train a graph neural network (GNN) model using semi-supervised learning. We evaluate this method on two large-scale datasets containing tweets related to climate change from June 2021 to June 2022 and gun control from January 2022 to January 2023. Our experiments demonstrate that enriching text-based embeddings of users with network information from the user interaction graph using our semi-supervised GNN method outperforms both classifiers trained on user textual embeddings and zero-shot classification using LLMs such as GPT4. We discuss the need for integrating nuanced understanding from social science with the scalability of computational methods to better understand how polarization on social media occurs for divisive issues such as climate change and gun control.
@@ I Introduction
Stance labeling remains a particularly challenging task due to the informal and noisy nature of social media language, the presence of sarcasm and irony, and the potential biases introduced by human annotators and automated models. To further facilitate research into polarization online, we develop a stance labeling method that incorporates both textual content and user social interactions in a scalable two-stage pipeline. In the first stage, we apply a reciprocal label propagation algorithm to the user-hashtag bipartite graph that generates a soft-labeled set of users. In the second stage, we construct a signed, weighted, and attributed user-user interaction network and use transformer language models to represent the content of user posts and to determine the sentiment of interactions between users. Using the soft labels generated from the label propagation stage, we employ semi-supervised training to train a graph neural network classifier to classify user stances.
Our stance labeling approach utilizes an understanding about linguistic and network homophily. Linguistic homophily states that people who share similar ideologies also share similar textual content, not only through similar keyword and hashtag usages but also through sentiment and linguistic features [8, 9], and the inverse phenomenon—that users with dissimilar views use different language—has also been investigated by researchers [2]. Likewise, the concept of network homophily, which describes the phenomenon that similar nodes tend to be connected to one another has long been studied in graph theory [10, 11] and informs the success of graph neural networks [12, 13, 14]. By combining the content-based and network-based approaches into a single, scalable pipeline for stance labeling, we facilitate further large-scale analysis of both interactional and affective polarization online.
@@
  • We develop an approach for estimating user stance labels that incorporates both textual features and social network interactions.
  • A two-stage pipeline allows our approach to scale to large datasets with soft labeling and semi-supervised training.
  • We discuss challenges and potential biases in stance determination of users when relying on short text content as commonly found on social media.
@@ II Related Work
As social media has become the primary forum for discussion on many topics, including political issues, there has been growing interest in user stance detection [15, 16, 17]. Many existing methods can be divided into two broad categories: textual-based methods that use the content of messages, and network-based methods that use social or conversational structure. Textual-based methods use the content of tweets, user profiles, and hashtags to determine a user’s stance on a particular topic. Several studies have used supervised classification of user stance labels where classifiers are trained using a set of features encoding tweet text, hashtags, and user profile information. These techniques have been applied to determine user political ideology [18, 19]. These approaches all generally rely on manually annotated sets of users to train the classifier models, and such methods are often limited by the size and diversity of the training content as manual annotation is both time-consuming and expensive. Furthermore, these methods can be susceptible to annotator bias and poor generalizability due to the limited distribution of training examples.
@@
Topic modeling and user clustering are two unsupervised techniques that have been employed for stance labeling. These unsupervised methods generally embed text or user profile information in a latent feature space and then apply a variety of topic modeling or cluster algorithms to identify texts of similar topics or clusters of similar users. While such techniques avoid the issues with manual annotation for supervised training, they suffer from the  curse of dimensionality—the search space for a solution grows exponentially with the increase in dimensionality—as there are many more possible patterns than in lower-dimensional subspaces. Selecting useful and informative features, and the computation and memory complexity of clustering methods on high-dimensional create problems for unsupervised clustering-based methods. For high-dimensional data, many clustering techniques fail to produce meaningful clusters, but on the other hand, such techniques can be very efficient on low-dimensional features spaces [16]. Several studies have therefore employed dimensionality reduction techniques prior to clustering to first project content embeddings to a low-dimensional space, and such techniques can be robust to class imbalance between different stance groups.
@@
Network-based approaches use the structure of social interaction networks to determine user ideology or stance. Such methods employ techniques such as label propagation, min-cut, and node embedding techniques to leverage structural information from retweet [17] and user follower [20] networks. Xiao et al. [21] generate a multi-relational network using retweets, mentions, likes, and follows for binary classification of user ideologies.
@@
Recent works have combined both textual content and network features in socially-infused text mining. Li and Goldwasser [22] combine user interactions and user sharing of news media to detect bias in news articles, and Johnson, Jin, and Goldwasser [23] combine multiple representations from lexical features and social interactions to categorize tweets by politicians on a variety of political topics. Pan et al. [24] use the network structure and node content to learn node representations for classifying the category of scientific publications. Such works combining textual and network information have demonstrated improvements in user stance deterction compared with methods that rely solely on textual content or network structure alone.
@@
Our approach is likewise informed by socially-infused text mining that combines information from text and network relationships for user stance labeling, but the approach we present in this paper is unique from the methods described above in that we i) combine language features from state-of-the-art transformer language models with the structure and sentiment of social interactions using graph neural networks, and ii) our two-stage process allows our method to scale to large datasets with minimal supervision.
@@ III Methods
This section describes our proposed method to estimate the stance of Twitter users on political issues. We first describe the collection of data for the gun control and climate changes datasets. Then we characterize the hashtag-based heuristic method to generate soft labels for two stance groups of users, which are used as seed users for training stance labeling methods. We then introduce several baseline methods followed by our GNN-based approach for stance labeling.
@@ III-A Data Collection
We collect tweets related to two prominent and contentious political issues—climate change and gun control. The keywords listed in Table I are used to scrape the initial set of relevant tweets related to each topic. To fully capture conversation cascades and user interactions, we then recursively collect all referenced tweets from each tweet in the initially matched set. This ensures that all available tweets from relevant conversations are included in our analysis, even if particular tweets do not explicitly use any of the keywords. The climate change dataset consists of tweets published between June 1st, 2021 and May 31st, 2022. In all, the climate change dataset consists of a total of 46M tweets authored by 4.8M unique users and contains 726,378 conversation threads of at least three tweets. The gun control dataset consists of tweets published between January 1st, 2022 and December 31st, 2022. In all, the gun control dataset consists of a total of 14.4M tweets from 2.66M unique users containing 335,000 conversation threads of at least three tweets.
@@
These issues are good models for understanding the differences in behavior on social media for topics that are highly event driven, like gun control, and those that are less focused on particular events, like climate change. Figure 1 illustrates that conversation threads related to climate change are started at a fairly consistent rate over time with some spikes in conversation activity connected with specific events. But compared with the conversation activity on Twitter surrounding gun control, illustrated in Figure 2, the frequency of discussions surrounding gun control spikes sharply in response to mass shooting events. The summer of 2022 saw several deadly and highly publicized mass shootings in Buffalo, Uvalde, and Chicago that generated large amounts of discourse surrounding gun control and gun violence on social media, illustrated by the sharp spikes from the end of May through late July. Shortly after mass shooting events though, the level of discourse abates, returning to near baseline levels between early August until November, when a shooting event at UVA caused a small spike in conversation activity. Event driven issues such as gun control exhibit different posting behavior and user interactions that we hypothesize lead to differences in polarization and require different techniques for intervention and mitigation of polarization on social media. Developing approaches for stance labeling that are both topic independent and robust to different types of discourse online is an important step for facilitating further analysis of polarization online.
Figure 1: Timeline of the number of Twitter conversations about climate change started each day.
Figure 2: Timeline of the number of Twitter conversations about gun control started each day.
@@ III-B User-Hashtag Stance Labeling
Based on a small (3-5) set of seed hashtags associated with each stance, we first propagate the stance associated with each seed hashtag across the user-hashtag bipartite graph to the set of users who posted a given labeled hashtag. The importance of a hashtag’s stance to a particular destination user is weighted by the number of times the user posted each hashtag, as described in Algorithm 2. Users are then assigned a label based on the number of hashtag usages for each stance group. User labels are then propagated back to the hashtags used by each user, as described in Algorithm 3. Each hashtag is then scored based on the difference between the normalized count of usages from users of each stance group. Hashtag scores are normalized via min-max normalization, and the mean and standard deviation of hashtag scores is computed. Hashtags that score one or more standard deviations above or below the mean are assigned the corresponding stance label. We then repeat alternating label propagation until the model converges, i.e. the set of labeled users is the same from the prior iteration to the next.
@@
While the user-hashtag bipartite label propagation method can quickly label a number of users with high precision. It suffers from the limitation that a large proportion of users do not use hashtags, and while it is well suited for determining hashtags that are highly associated with one particular stance group as compared with the other, hashtags that are used across both stances albeit in different contexts will remain unlabeled.
@@ III-C GNN Stance Labeling
The user-hashtag stance labeling method can efficiently determine the stance group of users and discover important hashtags associated with each stance group, but it is limited by the fact that not all users post hashtags and that hashtags consist of only a small part of the overall textual content contained in tweets. Furthermore, the method does not take into account any social interactions between users. The hashtag-based heuristic has high precision but relatively low recall due to the limited breadth of users that stance labeling method covers. To develop a scalable model capable of labeling users in large-scale datasets, we employ the user-hashtag bipartite method as a means to generate a soft-labeled set of users from which we train a graph neural network classifier over the attributed user-user interaction graph.
We define a user interaction as a retweet, mention, reply, or quote from the author of the tweet towards another user. We then construct a weighted, signed, and attributed user-user interaction network $G=(V,E,X)$, where $V$ is the set of users and a directed edge $(u,v)_{i}\in E$ connects the author user to the target user of their interaction as defined above. The edge weight $w(u,v)_{i}$ corresponds to the sentiment of the tweet, scored by a BERTweet-based sentiment analysis model. BERTweet is a RoBERTa model fine-tuned on English tweets and trained on the SemEval 2017 corpus [25]. To characterize the overall interactions between users, we consolidate the multiple edges between pairs of users corresponding to individual tweets by combining each directed edge $(u,v)_{i}$ into a single edge for each user pair $(u,v)$ with the composite edge weight $w(u,v)$ computed as the mean sentiment score across all tweets between two users. Each user $u\in V$ is also associated with a feature vector $x_{u}\in X$ which corresponds to the element-wise mean of the BERTweet embeddings of all tweets authored by the user. We train a graph neural network model as a stance label classifier using semi-supervised learning with the graph partially labeled with the set of seed users labeled by our hashtag-based heuristic as a training set.
For the GNN, we compare two implementations from the prominent families of graph neural networks—convolutional and attentional. For convolutional GNNs, the neighbor coefficients are fixed weights; whereas, in attentional GNNs these weights are computed implicitly by an attention mechanism. Specifically, we evaluate graphSAGE[13] as an example of a convolutional GNN and graph attention network (GAT) [26] as an example attentional GNN.
@@
where $\mathcal{\tilde{N}}(v)$ is the local neighborhood of node $v$ with added self-loop, $\boldsymbol{\mathrm{W}}$ is a trainable weight matrix, and $\sigma$ is a non-linear activation function. We use the mean aggregator to aggregate the neighboring node features.
We also consider a GAT layer operator, defined as:
@@
III-D Baselines
@@
We compare our stance labeling method against six transformer language models that leverage the textual content of tweets to represent users. With each language model, we compute a user embedding by computing the element-wise mean of the embedding of all the users’ tweets from BERT, RoBERTa, and BERTweet models. Using the set of labeled users from our heuristic method, we train an MLP classifier to predict user stance label.
We evaluate classification performance of the stance labeling methods using a set of manually annotated users and a set of users annotated via GPT-4 using zero-shot classification. Two social science experts independently annotated a set of 250 users for the climate change dataset and 350 users for the gun control dataset by analyzing the top 20 most popular tweets from each user. We provide the same set of tweets to GPT-4 and prompt the LLM to provide the sentiment of content with respect to gun control or climate change. For the gun control dataset, the function call prompt is: “The stance of the content. Is the message pro gun control, anti gun control, or neutral?” with three classification options of “pro”, “anti”, or “neutral”. For the climate change dataset, “The stance of the content. Does the message capture the user’s belief in climate change, disbelief, or is it neutral?” with classification options of “belief”, “disbelief”, or “neutral”. We then aggregate the stance labeling results for each individual tweet to determine a user’s overall stance on the issue of gun control or climate change. Users who had a majority of tweets classified as belonging to one stance were labeling accordingly and those users whose tweets were labeled as all neutral are classified as undetermined. In the case where there is an equivalent number of tweets labeled as both stances, the user is labeled as undetermined.
Using the set of manual and GPT-4 annotated users for the two datasets, we compare the classification performance of both text-based methods and our GNN-based method informed by socially infused text mining. We additionally compare the performance of GPT-4 for zero-shot stance labeling of users compared against the manual annotation by domain experts.